# Repetita Iuvant: Data Repetition Allows SGD to Learn High-Dimensional Multi-Index Functions

### v1

@@ 3.1 Main results
$y=h^{\star}(\langle\bm{w^{\star}},\bm{z}\rangle),\quad\text{with}\quad f(\bm{z},\bm{w},a)=a\sigma(\langle\bm{w},\bm{z}\rangle).$
@@
$\bm{w}_{t+1}=\frac{\bm{w}_{t}-\gamma\nabla^{\bot}_{\bm{w}}\mathcal{L}\left(f(\bm{z}^{\nu};\tilde{\bm{w}}_{t}(\rho),a_{0}),y^{\nu}\right)}{\left\lVert\bm{w}_{t}-\gamma\nabla^{\bot}_{\bm{w}}\mathcal{L}\left(f(\bm{z}^{\nu};\tilde{\bm{w}}_{t}(\rho),a_{0}),y^{\nu}\right)\right\rVert}$ (11)
where $\nabla^{\bot}f=\nabla f-\langle\nabla f,\bm{w}\rangle\bm{w}$ is the spherical gradient.
@@
The activation $\sigma$ is analytic, with bounded first and second derivative. Further, for all $n\geq 1$, we have
$\mathbb{E}\left[\sigma^{(n)}(x)\sigma^{\prime}(x)^{n-1}\right]\neq 0\quad\text{and}\quad\mathbb{E}\left[x\sigma^{(n)}(x)\sigma^{\prime}(x)^{n-1}\right]\neq 0$
Assumption 3.
The initial value $\bm{w}_{0}$ is drawn according to the uniform measure on
$\mathbb{S}_{\epsilon}=\{\bm{w}\in\mathbb{S}^{d-1}:\>\mathrm{sign}(\langle\bm{w},\bm{w^{\star}}\rangle)=\epsilon\}$
where $\epsilon$ depends on $h^{\star}$ and $\sigma$.
@@
Suppose that Assumptions 1 and 2 hold, and let $\rho=\rho_{0}d^{-1}$. Then, for any $\delta>0$, there exists constants $\gamma_{0}(\delta)$ and $C(\delta)$ such that the following holds:
@@
The above holds for almost every choice of $\rho_{0}$ under the Lebesgue measure.
@@
We provide here the proof sketch for Theorem 2. For any vector $\bm{w}$, under the correlation loss, we have
$\nabla_{\bm{w}}\mathcal{L}(f(\bm{z};\bm{w},a_{0}),y)=a_{0}y\sigma^{\prime}(\langle\bm{w},z\rangle)\cdot\bm{z},$ (12)
so the gradient is aligned with $\bm{z}$. As a result, we have
$\nabla_{\bm{w}}\mathcal{L}\left(f(\bm{z};\tilde{\bm{w}}(\rho),a_{0}),y\right)=a_{0}y\sigma^{\prime}\big{(}\langle\bm{w},\bm{z}\rangle+a_{0}\rho y\sigma^{\prime}(\langle\bm{w},\bm{z}\rangle)\cdot\underbrace{\left\lVert\bm{z}\right\rVert^{2}}_{\approx d}\big{)}\cdot\bm{z}$ (13)
@@
  • Even though $\rho\asymp d^{-1}$, since the gradient lies along $\bm{z}$, the additional dot product with $\bm{z}$ amplifies the signal by a factor of $d$,
  • The resulting gradient (13) is a non-linear function of $y$, as opposed to the linear function of (12).
@@ 3.2 Discussion
Assumption 1 simplifies Alg. 1 to a more tractable version. In particular, the normalization step allows us to keep track of only one parameter, the overlap between $\bm{w}_{t}$ and $\bm{w}^{\star}$. We expect however Thm.2 to hold in the general setting of Alg.1.
Assumption 2 is a regularity assumption, ensuring that some quantities of interest in the proof are non-zero. The first part is satisfied by virtually every activation function except for ReLU; the second is more restrictive, but we show that it is satisfied for biased activations:
@@
Assumption 3 is similar to the one in Ben Arous et al. (2021), and is necessary in case the initialization is a strict saddle. It is equivalent to conditioning the usual uniform initialization on an event of probability $1/2$.
@@
A succinct summary of the proof goes as follows: the first time we see a sample $\bm{z}$, we store information about $y$ along the direction of $\bm{z}$. The second time we see the same sample, the component along $\bm{z}$ in $\bm{w}$ interact non-linearly with $y$, which bypasses the CSQ framework. Importantly, we expect that this behavior also appears in SGD algorithms where we do not see the data twice in a row, but across multiple epochs! Indeed, upon small enough correlation between the $\bm{z}^{\nu}$, seeing multiple examples before the repetition of $\bm{z}$ should not interfere with the data stored along this direction. We are therefore not claiming that Algorithm 1 is superior to other classically-used methods; rather, it provides a tractable and realistic setting to study data repetition in SGD, which was ignored in previous theoretical works on the topic.
(a) $h^{\star}(x)=\operatorname{relu}(x)$
(b) $h^{\star}(x)=\mathrm{He}_{2}(x)$
(c) $h^{\star}(x)=\mathrm{He}_{3}(x)$
(d) $h^{\star}(x)=\mathrm{He}_{4}(x)$
Figure 1: Learning single-index targets – Evolution of the Cosine Similarity attained by EgD (crosses) and SGD (dots) as a function of normalized iteration time. The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. (a) $(\ell,\ell^{\star})=(1,1)$: both the algorithm learn in linear time. (b) $(\ell,\ell^{\star})=(2,2)$: both the algorithm learn in $O(d\log d)$ time. (c) $(\ell,\ell^{\star})=(3,1)$: EgD learns in linear time, while SGD would learn the same function in quadratic time. (d) $(\ell,\ell^{\star})=(4,2)$: EgD learns in $O(d\log d)$ time, while SGD would learn in $O(d^{3})$ time. See all the details in App. C.
@@ 3.3 Numerical illustrations
We illustrate in Fig. 1 the stark difference between the weak recovery dynamics of One-Pass SGD and Algo. 1 for single-index targets. The Cosine Similarity between the learned weight $\bm{w}_{t}$ and the ground truth informative direction $\bm{w}_{\star}$ is shown as a function of the time steps $t$ for different target functions. Two optimization routines are considered to exemplify the learning behaviour: a) Extragradient Descent (EgD), corresponding to the family of Algorithms 1 with positive $\rho$ parameters; b) One-Pass SGD, corresponding to vanilla SGD, or equivalently Algorithm 1 associated with $\rho=0$ hyperparameter. The scaling of $(\gamma,\rho)$ as a function of the input dimension $d$ and exponent $\ell^{\star}$ are given in Thm. 2, while the mini-batch size is fixed to $n_{b}=1$ (See App. B for extension to larger $n_{b}$). For all plots, we take $\sigma$ to be the $\mathrm{relu}$ activation and the implementation details can be found in Appendix C.
@@ 4 Multi-Index Model
We now investigate the superiority of Algorithm 1 over the idealized One-Pass SGD training scheme when learning multi-index targets. In this setting, the networks weights can align with only a few select directions from $V^{\star}$. To quantify this behavior, we define directional versions of the Information and Generative Exponents:
@@
Let $\bm{v}\in V^{\star}$. The Information Exponent $\ell_{\bm{v}}$ (resp. Generative exponent $\ell^{\star}_{\bm{v}}$) of $y$ in the direction $\bm{v}$ is the smallest $k$ such that
@@
$\mathbb{E}\left[yH_{k}(\langle\bm{v},\bm{z}\rangle)\right]\neq 0\quad(\text{resp.}\ \mathbb{E}\left[f_{\bm{v}}(y)H_{k}(\langle\bm{v},\bm{z}\rangle)\right]\neq 0)$
for a function $f_{\bm{v}}$. We also let $\ell=\min_{\bm{v}}\ell_{\bm{v}}$ and $\ell^{\star}=\min_{\bm{v}}\ell^{\star}_{\bm{v}}$.
@@
The identification of the class of hard functions has been subject of intense theoretical scrutiny (Abbe et al., 2021, 2022, 2023; Dandi et al., 2023; Bietti et al., 2023); and for the problem of initial alignment it was shown to depend on the Information exponent defined above. Thus, similarly to the single-index scenario, the time complexity needed for Algorithm 1 to weakly recover the target subspace follows the law in eq. (8) for the multi-index information exponent $\ell$. A natural question is thus to ask whether the class of Algorithm 1 also allows us to bypass this requirement.
Our simulations enable us to affirmatively answer this question. We illustrate this numerically by comparing the performance of One-Pass SGD with Algorithm 1 in Figure 2; we show that the dynamics of the latter algorithm is governed by the multi-index Generative Exponent (Def. 3).
(a) $h^{\star}(x_{1},x_{2})=x_{1}+x_{1}x_{2}$
(b) $h^{\star}(x_{1},x_{2})=\mathrm{sign}{(x_{1}x_{2})}$
(c) $h^{\star}(x_{1},x_{2})=x_{1}+\mathrm{He}_{3}(x_{2})$
(d) $h^{\star}(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$
Figure 2: Learning multi-index targets – Evolution of the maximum Cosine Similarities attained by EgD (crosses) and SGD (dots) as a function of the normalized iteration time. Different target directions $\{\bm{w}^{\star}_{r}\}_{r\in[k]}$ are identified by different colors: $\bm{w}^{\star}_{1}$ (blue), $\bm{w}^{\star}_{2}$ (orange), $\bm{w}^{\star}_{3}$ (green). The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. See details in App. C (a) $(\ell,\ell^{\star})=(1,1)$: both the algorithm learn the first direction in linear time, and they also learn the second in another $O(d)$ steps using the staircase mechanism. (b) $(\ell,\ell^{\star})=(2,2)$: both the algorithm learn the two direction simultaniously in $O(d\log d)$ steps. (c) $(\ell_{\bm{w}^{\star}_{2}},\ell_{\bm{w}^{\star}_{2}}^{\star})=(3,1)$: both the algorithms learn $\bm{w}^{\star}_{1}$ in linear time, but only EgD can also learn $\bm{w}^{\star}_{2}$ in $O(d)$; SGD requires $O(d^{2})$ steps to learn $\bm{w}^{\star}_{2}$. (d) $(\ell,\ell^{\star})=(3,2)$: EgD learns all the 3 directions simultaneously in $O(d\log d)$ steps, while SGD needs quadratic time to achieve the same recovery.
4.1 Numerical investigation
@@
We measure the maximal cosine similarity between the hidden layer neurons and the different rows of the target’s weight matrix $W^{\star}$ as a function of the iteration time. We set without loss of generality the target’s directions to the standard basis of $\mathbb{R}^{d}$ to ease the notation. The algorithms considered are again EgD and vanilla One-Pass SGD. We again refer to Appendix C for the implementation details.
@@
SGD-easy non-symmetric targets ($\ell\!=\!\ell^{\star}\!=\!1$) – Fig. 2, left, shows $h^{\star}(x_{1},x_{2})=x_{1}+x_{1}x_{2}$. Here both SGD and Alg. 1 learn in $T=O(d)$ the directions $(\bm{w}^{\star}_{1},\bm{w}^{\star}_{2})$.
@@
SGD-easy symmetric targets ($\ell\!=\!\ell^{\star}\!=\!2$) – Fig. 2, center left, shows $h^{\star}(x_{1},x_{2})=x_{1}x_{2}$. Here both SGD and Alg. 1 learn in $T=O(d\log d)$ the directions $(\bm{w}^{\star}_{1},\bm{w}^{\star}_{2})$.
@@
Non-symmetric targets with SGD-easy and hard directions ($\ell_{\bm{v}}>2,\ell^{\star}_{\bm{v}}=1$) – Fig. 2, center-right, shows $h^{\star}(x_{1},x_{2})\!=\!x_{1}\!+\!\mathrm{He}_{3}(x_{2})$. Here both SGD and Alg. 1 learn in $T=O(d)$ the direction $\bm{w}^{\star}_{1}$, that satisfies $\ell_{\bm{w_{1}^{\star}}}=\ell_{\bm{w_{1}}^{\star}}^{\star}=1$. However, since $\ell_{\bm{w_{2}^{\star}}}=3$ while $\ell_{\bm{w_{2}^{\star}}}^{\star}=1$, One-Pass SGD suffers from the limitations detailed in eq. (8) and requires $\Omega(d^{2})$ samples to learn the second direction $\bm{w}^{\star}_{2}$, . This contrasts with the behaviour of Alg. 1 that learns also $\bm{w}^{\star}_{2}$ in $T=O(d)$ steps.
@@
SGD-hard symmetric targets ($\ell>2,\ell^{\star}=2$) – The right section Fig. 2 shows $h^{\star}(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$. The Information Exponent is $\ell=3$ and hence vanilla SGD is not able to learn any direction in the target subspace in $T=O(d\log d)$ iterations. However, Algorithms 1 learns in $T=O(d\log d)$ steps all the three directions $\{\bm{w}^{\star}_{1},\bm{w}^{\star}_{2},\bm{w}^{\star}_{3}\}$ since the Generative Exponents $\ell_{\bm{w}^{\star}_{r}}^{\star}$ are all equal to two.

### v2

@@ 3.1 Main results
$y=h^{\star}(\langle\mathbf{w^{\star}},\mathbf{z}\rangle),\quad\text{with}\quad f(\mathbf{z},\mathbf{w},a)=a\sigma(\langle\mathbf{w},\mathbf{z}\rangle).$
@@
$\mathbf{w}_{t+1}=\frac{\mathbf{w}_{t}-\gamma\nabla^{\bot}_{\mathbf{w}}\mathcal{L}\left(f(\mathbf{z}^{\nu};\tilde{\mathbf{w}}_{t}(\rho),a_{0}),y^{\nu}\right)}{\left\lVert\mathbf{w}_{t}-\gamma\nabla^{\bot}_{\mathbf{w}}\mathcal{L}\left(f(\mathbf{z}^{\nu};\tilde{\mathbf{w}}_{t}(\rho),a_{0}),y^{\nu}\right)\right\rVert}$ (12)
where $\nabla^{\bot}f=\nabla f-\langle\nabla f,\mathbf{w}\rangle\mathbf{w}$ is the spherical gradient.
@@
The activation $\sigma$ is analytic almost everywhere, and $\sigma^{(n)}$ satisfies the polynomial growth assumption (3) for $n\in\mathbb{N}$ (possibly for different constants $C_{n}$). Further, for all $n\geq 1$, we have
$\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[\sigma^{(n)}(x)\sigma^{\prime}(x)^{n-1}\right]\neq 0\quad\text{and}\quad\mathbb{E}_{x\sim\mathcal{N}(0,1)}\left[x\sigma^{(n)}(x)\sigma^{\prime}(x)^{n-1}\right]\neq 0$
@@
Suppose that Assumptions 1-4 hold, and let $\rho=\rho_{0}d^{-1}$. Then, under an event of probability $1/2$ on the initialization $(a_{0},\mathbf{w}_{0})$, for each $\delta>0$, there exist constants $\gamma_{0}(\delta)$ and $C(\delta)$ such that the following holds:
@@
The above holds for almost every choice of $\rho_{0}$ under the Lebesgue measure. In particular, in a network with $p=O(\log\frac{1}{\delta})$ neurons, at least one of them achieves weak recovery with probability $1-\delta$.
@@
We provide here the proof sketch for Theorem 2. For any vector $\mathbf{w}$, under the correlation loss, we have
$\nabla_{\mathbf{w}}\mathcal{L}(f(\mathbf{z};\mathbf{w},a_{0}),y)=a_{0}y\sigma^{\prime}(\langle\mathbf{w},z\rangle)\cdot\mathbf{z},$ (13)
so the gradient is aligned with $\mathbf{z}$. As a result, we have
$\nabla_{\mathbf{w}}\mathcal{L}\left(f(\mathbf{z};\tilde{\mathbf{w}}(\rho),a_{0}),y\right)=a_{0}y\sigma^{\prime}\big{(}\langle\mathbf{w},\mathbf{z}\rangle+a_{0}\rho y\sigma^{\prime}(\langle\mathbf{w},\mathbf{z}\rangle)\cdot\underbrace{\left\lVert\mathbf{z}\right\rVert^{2}}_{\approx d}\big{)}\cdot\mathbf{z}$ (14)
@@
  • Even though $\rho\asymp d^{-1}$, since the gradient lies along $\mathbf{z}$, the additional dot product with $\mathbf{z}$ amplifies the signal by a factor of $d$,
  • The resulting gradient (14) is a non-linear function of $y$, as opposed to the linear one of (13).
@@ 3.2 Discussion
Assumption 3 simplifies Alg. 1 to a more tractable version. In particular, the normalization step allows us to keep track of only one parameter, the overlap between $\mathbf{w}_{t}$ and $\mathbf{w}^{\star}$. We expect however Thm. 2 to hold in the general setting of Alg. 1.
Assumption 4 is a regularity assumption, ensuring that some quantities of interest in the proof are non-zero. The first part is satisfied by virtually every activation function; the second is more restrictive, but we show that it is satisfied for biased activations:
@@
A succinct summary of the proof goes as follows: the first time we see a sample $\mathbf{z}$, we store information about $y$ along the direction of $\mathbf{z}$. The second time we see the same sample, the component along $\mathbf{z}$ in $\mathbf{w}$ interact non-linearly with $y$, which bypasses the CSQ framework. Importantly, we expect that this behavior also appears in SGD algorithms where we do not see the data twice in a row, but across multiple epochs! Indeed, upon small enough correlation between the $\mathbf{z}^{\nu}$, seeing multiple examples before the repetition of $\mathbf{z}$ should not interfere with the data stored along this direction. We are therefore not claiming that Algorithm 1 is superior to other classically-used methods; rather, it provides a tractable and realistic setting to study data repetition in SGD, which was ignored in previous theoretical works on the topic.
(a) $\scriptstyle h^{\star}(x)=\operatorname{relu}(x)$
(b) $\scriptstyle h^{\star}(x)=\mathrm{He}_{2}(x)$
(c) $\scriptstyle h^{\star}(x)=\mathrm{He}_{3}(x)$
(d) $\scriptstyle h^{\star}(x)=\mathrm{He}_{4}(x)$
Figure 1: Learning single-index targets – Evolution of the Cosine Similarity attained by EgD (crosses) and SGD (dots) as a function of the normalized iteration time. (dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance). (a) $(\ell,\ell^{\star})=(1,1)$: both algorithms learn in linear time. (b) $(\ell,\ell^{\star})=(2,2)$: both algorithms learn in $O(d\log d)$ time. (c) $(\ell,\ell^{\star})=(3,1)$: EgD learns in linear time, while SGD requires $O(d^{2})$ time. (d) $(\ell,\ell^{\star})=(4,2)$: EgD learns in $O(d\log d)$ time, SGD requires $O(d^{3})$ time (Details in App. E).
@@ 3.3 Numerical illustrations
We illustrate in Fig. 1 the stark difference between the weak recovery dynamics of One-Pass SGD and Algo. 1 for single-index targets. The Cosine Similarity between the learned weight $\mathbf{w}_{t}$ and the ground truth informative direction $\mathbf{w}_{\star}$ is shown as a function of the time steps $t$ for different target functions. Two optimization routines are considered to exemplify the learning behaviour: a) Extragradient Descent (EgD), corresponding to the family of Algorithms 1 with positive $\rho$ parameters; b) One-Pass SGD, corresponding to vanilla SGD, or equivalently Algorithm 1 associated with $\rho=0$ hyperparameter. The scaling of $(\gamma,\rho)$ as a function of the input dimension $d$ and exponent $\ell^{\star}$ are given in Thm. 2, while the mini-batch size is fixed to $n_{b}=1$ (See App. D for extension to larger $n_{b}$). For all plots, we take $\sigma=\mathrm{relu}$ and we refer to App. E for details.
@@
(a) $\scriptstyle h^{\star}(\mathbf{x})=x_{1}+x_{1}x_{2}$
(b) $\scriptstyle h^{\star}(\mathbf{x})=\mathrm{sign}{(x_{1}x_{2})}$
(c) $\scriptstyle h^{\star}(\mathbf{x})=x_{1}+\mathrm{He}_{3}(x_{2})$
(d) $\scriptstyle h^{\star}(\mathbf{x})=x_{1}x_{2}x_{3}$
Figure 2: Learning multi-index targets – Evolution of the maximum Cosine Similarities attained by EgD (crosses) and SGD (dots) as a function of the normalized iteration time. The Different directions $\{\mathbf{w}^{\star}_{r}\}_{r\in[k]}$ are identified by colors: $\mathbf{w}^{\star}_{1}$ (blue), $\mathbf{w}^{\star}_{2}$ (orange), $\mathbf{w}^{\star}_{3}$ (green). (dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance). (a) $(\ell,\ell^{\star})=(1,1)$: both algorithms learn the first direction in linear time, as well the second one in $O(d)$ steps using the staircase mechanism. (b) $(\ell,\ell^{\star})=(2,2)$: both algorithms learn the two directions simultaneously in $O(d\log d)$ steps. (c) $(\ell_{\mathbf{w}^{\star}_{2}},\ell_{\mathbf{w}^{\star}_{2}}^{\star})=(3,1)$: both algorithms learn $\mathbf{w}^{\star}_{1}$ in linear time, but only EgD learns $\mathbf{w}^{\star}_{2}$ in $O(d)$; SGD instead requires $O(d^{2})$. (d) $(\ell,\ell^{\star})=(3,2)$: EgD learns all 3 directions simultaneously in $O(d\log d)$ steps, while SGD needs $O(d^{2})$ time. (Details in App. E).
@@ 4 Multi-Index Model
We now investigate the superiority of Algorithm 1 over the idealized One-Pass SGD training scheme when learning multi-index targets. In this setting, the networks weights can align with only a few select directions from $V^{\star}$. To quantify this behavior, we define directional versions of the Information and Generative Exponents (see also (Troiani et al., 2024)):
@@
Let $\mathbf{v}\in V^{\star}$. The Information Exponent $\ell_{\mathbf{v}}$ (resp. Polynomial Generative Exponent $\ell^{\star}_{\mathbf{v}}$) of $y$ in the direction $\mathbf{v}$ is the smallest $k$ such that
@@
$\mathbb{E}\left[yH_{k}(\langle\mathbf{v},\mathbf{z}\rangle)\right]\neq 0\quad(\text{resp.}\ \mathbb{E}\left[p_{\mathbf{v}}(y)H_{k}(\langle\mathbf{v},\mathbf{z}\rangle)\right]\neq 0)$
for a polynomial $p_{\mathbf{v}}$. We also let $\ell=\min_{\mathbf{v}}\ell_{\mathbf{v}}$ and $\ell_{p}^{\star}=\min_{\mathbf{v}}\ell^{\star}_{\mathbf{v}}$. The minimal subspace $U^{\star}\subseteq V^{\star}$ is defined as the smallest subspace $U$ such that $\ell_{\mathbf{v}}>\ell$ (resp. $\ell_{\mathbf{v}}^{\star}>\ell_{p}^{\star}$) for every $v\in U^{\bot}$.
@@
The identification of the class of hard functions has been subject of intense theoretical scrutiny (Abbe et al., 2021, 2022, 2023; Dandi et al., 2024a; Bietti et al., 2023; Troiani et al., 2024); and for the problem of initial alignment it was shown to depend on the Information exponent defined above. Thus, similarly to the single-index scenario, the time complexity needed for Algorithm 1 to weakly recover the target subspace follows the law in eq. (9) for the multi-index information exponent $\ell$. A natural question is thus to ask whether the class of Algorithm 1 also allows us to bypass this requirement. We show that this is indeed the case. Consider a neural network as in Equation (1); we assume that the internal learning rate $\rho$ is chosen independently for each neuron:
Assumption 5.
For each $j\in[p]$, the parameter $\rho_{0,j}$ is chosen independently from all other neurons, according to a symmetric distribution which is absolutely continuous w.r.t the Lebesgue measure.
Theorem 4.
Let $\delta>0$, and $\Pi^{\star}$ the projection operator on $V^{\star}$. There exists constants $\gamma_{0}(\delta),C(\delta)$ depending only on $\delta$ such that if we take $p\geq C(\delta)k$ and
  • if $\ell_{p}^{\star}=1$, $\gamma=\gamma_{0}(\delta)d^{-1}$ and $T=C(\delta)d$,
  • if $\ell_{p}^{\star}=2$, $\gamma=\gamma_{0}(\delta)(d\log(d))^{-1}$ and $T=C(\delta)d\log(d)^{2}$,
then with probability $1-\delta$, there exists a $t\leq T$ such that at least a positive proportion of the vectors $\mathbf{w}_{t,i}$ satisfy $\left\lVert\Pi^{\star}\mathbf{w}_{t,i}\right\rVert\geq\eta$. Further, if $\ell_{p}^{\star}=1$, then the vectors $\{\Pi^{\star}\mathbf{w}_{t,i}\}_{i\in[p]}$ span the minimal subspace $U^{\star}$.
(a) ${\scriptstyle h^{\star}(\mathbf{x})=x_{1}+x_{1}\mathrm{He}_{3}(x_{2})}$
(b) ${\scriptstyle h_{\mathrm{sign}}^{\star}(\mathbf{x})=\mathrm{sign}(x_{1}x_{2}x_{3}),\quad h^{\star}_{\mathrm{stair}}(\mathbf{x})=\mathrm{He}_{2}(x_{1})+\mathrm{sign}(x_{1}x_{2}x_{3})}$
Figure 3: Hierarchical learning – Evolution of the maximum Cosine Similarities with different target directions $\{\mathbf{w}^{\star}_{r}\}_{r\in[k]}$ are identified by different colors: $\mathbf{w}^{\star}_{1}$ (blue), $\mathbf{w}^{\star}_{2}$ (orange), $\mathbf{w}^{\star}_{3}$ (green) as a function of the normalized iteration time. (a): both the algorithms learn $\mathbf{w}^{\star}_{1}$ in linear time, but only EgD can also learn $\mathbf{w}^{\star}_{2}$ in $O(d)$ via a staircase mechanism; SGD requires another $O(d^{2})$ steps to take advantage of the staircase and learn $\mathbf{w}^{\star}_{2}$. (b): EgD performance for the target function $f^{\star}_{\mathrm{stair}}$ (dots) and $f^{\star}_{\mathrm{sign}}$(squares); we plot in separate subplots the overlap with different directions to highlight the presence of hierarchical learning mechanism. Learning the first target direction $\mathbf{w}^{\star}_{1}$ triggers the hierarchical mechanism and EgD is able to weakly recover the full target subspace $\mathrm{Span}(\mathbf{w}^{\star}_{1},\mathbf{w}^{\star}_{2},\mathbf{w}^{\star}_{3})$ in $O(d\log d)$, while this does not happen when removing $\mathrm{He}_{2}(x_{1})$ from the target. See App. E for additional details.
4.1 Numerical investigation with multi-index target
@@
We illustrate the results of Thm. 4 numerically by comparing the performance of One-Pass SGD with Alg. 1 in Fig 2. We measure the maximal cosine similarity between the hidden layer neurons and the different rows of the target’s weight matrix $W^{\star}$ as a function of the iteration time. We set without loss of generality the target’s directions to the standard basis of $\mathbb{R}^{d}$ to ease the notation. The algorithms considered are again EgD and vanilla One-Pass SGD (See App. E for details).
@@
  • SGD-easy non-symmetric targets ($\ell\!=\!\ell^{\star}\!=\!1$) – Fig. 2, left, shows $h^{\star}(x_{1},x_{2})=x_{1}+x_{1}x_{2}$. Here both SGD and Alg. 1 learn in $T=O(d)$ the directions $(\mathbf{w}^{\star}_{1},\mathbf{w}^{\star}_{2})$.
@@
  • SGD-easy symmetric targets ($\ell\!=\!\ell^{\star}\!=\!2$) – Fig. 2, center left, shows $h^{\star}(x_{1},x_{2})=x_{1}x_{2}$. Here both SGD and Alg. 1 learn in $T=O(d\log d)$ the directions $(\mathbf{w}^{\star}_{1},\mathbf{w}^{\star}_{2})$.
@@
  • Non-symmetric targets with SGD-easy and hard directions ($\ell_{\mathbf{v}}>2,\ell^{\star}_{\mathbf{v}}=1$) – Fig. 2, center-right, shows $h^{\star}(x_{1},x_{2})\!=\!x_{1}\!+\!\mathrm{He}_{3}(x_{2})$. Here both SGD and Alg. 1 learn in $T=O(d)$ the direction $\mathbf{w}^{\star}_{1}$, that satisfies $\ell_{\mathbf{w_{1}^{\star}}}=\ell_{\mathbf{w_{1}}^{\star}}^{\star}=1$. However, since $\ell_{\mathbf{w_{2}^{\star}}}=3$ while $\ell_{\mathbf{w_{2}^{\star}}}^{\star}=1$, One-Pass SGD suffers from the limitations detailed in eq. (9) and requires $\Omega(d^{2})$ samples to learn the second direction $\mathbf{w}^{\star}_{2}$. This contrasts with the behaviour of Alg. 1 that learns also $\mathbf{w}^{\star}_{2}$ in $T=O(d)$ steps.
@@
  • SGD-hard symmetric targets ($\ell>2,\ell^{\star}=2$) – The right section Fig. 2 shows $h^{\star}(x_{1},x_{2},x_{3})=x_{1}x_{2}x_{3}$. The Information Exponent is $\ell=3$, hence vanilla SGD does not learn any direction in the target subspace in $T=O(d\log d)$ iterations. However, Algorithms 1 learns in $T=O(d\log d)$ steps all the three directions $\{\mathbf{w}^{\star}_{1},\mathbf{w}^{\star}_{2},\mathbf{w}^{\star}_{3}\}$ since the Generative Exponents $\ell_{\mathbf{w}^{\star}_{r}}^{\star}$ are all equal to two.